# Center vortices and localized Dirac modes in the deconfined phase of 2+1 dimensional lattice $\mathbb{Z}_{2}$ gauge theory

### v1

@@ Abstract
We study the deconfinement transition in 2+1 dimensional lattice $\mathbb{Z}_{2}$ gauge theory both as a percolation transition of center vortices and as a localization transition for the low-lying Dirac modes. We study in detail the critical properties of the Anderson transition in the Dirac spectrum in the deconfined phase, showing that it is of BKT type; and the critical properties of the center-vortex percolation transition, showing that they differ from those of ordinary two-dimensional percolation. We then study the relation between localized modes and center vortices in the deconfined phase, identifying the simple center-vortex structures that mainly support the localized Dirac modes. As the system transitions to the confined phase, center vortices merge together into an infinite cluster, causing the low Dirac modes to delocalize.
@@ I Introduction
A growing amount of evidence shows that the localization properties of low-lying Dirac modes in a gauge theory are strongly connected with its confining properties: Low Dirac modes, while extended in the confined phase of a gauge theory, become localized in the deconfined phase, exactly at the critical temperature when this corresponds to a genuine phase transition [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] (see Ref. [24] for a recent review). More precisely, in the deconfined phase the localized low Dirac modes are separated from the delocalized bulk modes by a critical point in the spectrum known as “mobility edge”. At the mobility edge the localization length diverges and a phase transition along the spectrum, known as Anderson transition [25, 26, 27, 28, 29], takes place. As one crosses over to the confined phase, the mobility edge vanishes (possibly discontinuously [17, 16]), and localized modes disappear from the spectrum. While the existence of a relation between localization and deconfinement is by now established, the nature of this relation has not been fully elucidated yet.
@@
The connection between deconfinement and localization of low Dirac modes in gauge theories is qualitatively understood in terms of the “sea/islands” picture proposed in Ref. [5] and further developed in Refs. [30, 31, 32, 24, 20]. In the “physical sector” where Polyakov loops order along 1, selected by dynamical fermions or, in the case of pure gauge theories, by (infinitely) heavy external fermionic probes, a pseudogap opens in the Dirac spectrum due to the ordering of the Polyakov loop. In the resulting “sea” of ordered local Polyakov loops, gauge-field fluctuations that reduce the correlation across time slices, such as Polyakov-loop fluctuations away from 1, are localized on “islands”, typically well-separated in space. These fluctuations reduce the ability of Dirac modes to diffuse in space, and can “trap” the Dirac modes if these develop an eigenvalue below the gap, localizing the corresponding eigenvectors. This mechanism clearly applies not only in a deconfined phase, but more generally when the Polyakov loop gets ordered, e.g., in the Higgs phase of the $\mathrm{SU}(2)$ Higgs model [22].
@@
The connection between localization and deconfinement shows up even in the simplest gauge theory with a deconfining transition, namely pure $\mathbb{Z}_{2}$ gauge theory in 2+1 dimensions [19]. This model is dual to an anisotropic three-dimensional Ising model [33] (summed over boundary conditions: see Refs. [34, 35] and references therein, and Ref. [20] for a detailed derivation), which allows for an accurate determination of the critical value of the gauge coupling [34] through the use of a cluster algorithm [36, 37]. While localization of low Dirac modes was clearly demonstrated in Ref. [19], a precise determination of the mobility edge and of its temperature dependence, as well as a detailed study of the corresponding Anderson transition, were not undertaken.
@@
Gauge configurations in an Abelian model like the $\mathbb{Z}_{2}$ theory are entirely determined by the simplest plaquettes and by the Polyakov loops in the various directions, with those along the spatial ones becoming irrelevant in the large-volume limit. In the $\mathbb{Z}_{2}$ theory, plaquette configurations on the direct lattice are in one-to-one correspondence with configurations of closed loops on the dual lattice, constructed by associating an “active” dual link to each nontrivial plaquette. Dual loops can further intersect each other, and are conveniently grouped in disconnected clusters of various sizes. Since the gauge group is Abelian and so equal to its own center, these clusters are trivially the same as the center vortices of this model, so we will refer to them also as center vortices in the following (for recent results on center vortices in QCD and a list of references, see Ref. [38]).
@@
Clearly, the sum of the sizes of these clusters equals the number of negative plaquettes of the gauge configuration on the direct lattice, and one can exactly recast the partition function in terms of dual clusters. The dynamics of $\mathbb{Z}_{2}$ gauge fields is then the dynamics of center vortices. In this language, confinement can be understood as center-vortex percolation, i.e., as the formation of a percolating cluster of dual active links, that leads to a nonvanishing string tension. Conversely, deconfinement corresponds to the disappearance of the percolating cluster, with only clusters of finite size surviving the large-volume limit. [^1] Notice that since a plaquette configuration is left unchanged by a center transformation, the center sector selected by the Polyakov loop, as it gets ordered and breaks center symmetry, plays no role in this dynamics. The idea of looking at the confinement/deconfinement transition as a center-vortex percolation transition was proposed in Ref. [39] for the pure $\mathrm{SU}(2)$ gauge theory. Reference [40] studied instead $\mathbb{Z}_{2}$ gauge theory in 2+1 dimensions, showing the presence of an infinite center-cluster in typical gauge configurations in the confined phase, and demonstrating its crucial role in establishing confinement. On the other hand, the behavior of clusters across the deconfinement transition was not investigated.
@@
Pure $\mathbb{Z}_{2}$ gauge theory in 2+1 dimensions is clearly the easiest setup in which one can study the nature of the gauge-field fluctuations relevant to localization, and how these relate to the confining properties of the theory. In particular, the connection between deconfinement and localization on the one hand, and that between confinement and percolation on the other, suggests that the features of the non-percolating center clusters expected in the deconfined phase should play an important role for localization. Identifying the features of gauge-field configurations that connect localization and deconfinement in this simple model may lead to further insight on the problem of confinement in the physically relevant case of QCD, and more generally in non-Abelian gauge theories.
@@
In this paper we investigate in detail the localization of Dirac modes, the percolation of center vortices, and how they affect each other in pure $\mathbb{Z}_{2}$ gauge theory in 2+1 dimensions on the lattice, probed with external staggered fermions. After briefly introducing the model, in Sec. II we discuss a few generalities about center vortices and localization. In Sec. III we present our numerical results. We first determine the nature of the Anderson transition and the temperature dependence of the mobility edge. We then study the percolation of center vortices across the deconfinement transition and the correlation between center clusters and localized modes, and test the refined sea/islands picture of Ref. [20]. We then identify the features of center vortices mostly relevant to localization. Finally, in Sec. IV we present our conclusions and discuss prospects for future studies.
@@ II Localization and center vortices in $\mathbb{Z}_{2}$ lattice gauge theory at finite temperature
In this Section, after briefly describing 2+1-dimensional pure $\mathbb{Z}_{2}$ gauge theory on the lattice and how to probe it with external staggered fermions, we summarize the relevant aspects of center vortices and their percolation, and of Dirac mode localization.
II.1 Finite-temperature $\mathbb{Z}_{2}$ lattice gauge theory in 2+1 dimensions
We consider pure $\mathbb{Z}_{2}$ gauge theory on a 2+1-dimensional hypercubic $N_{t}\times N_{s}^{2}$ lattice. Lattice sites are denoted by $n=(n_{1},n_{2},n_{3})=(t,\vec{x})$, where $10\leq t=n_{1}\leq N_{t}-$ 1, $10\leq x_{1,2}=n_{2,3}\leq N_{s}-$ 1. The dynamical variables of the model are the link variables $U_{\mu}(n)=\pm 1$, taking values in $\mathbb{Z}_{2}$, and associated with the oriented lattice links connecting a site $n$ with its neighbors $n+\hat{\mu}$, where $\mu=1,2,3$ and $\hat{\mu}$ is the unit lattice vector in direction $\mu$. Periodic boundary conditions are imposed in all directions.
@@ II.1 Finite-temperature $\mathbb{Z}_{2}$ lattice gauge theory in 2+1 dimensions
Together with the analogues of $P(\vec{x})$ winding around the spatial directions, these quantities determine the link configuration up to a gauge transformation.
@@ II.2 Staggered Dirac spectrum
We probe the gauge-field configurations using external fermion fields by studying the spectrum of the staggered Dirac operator, which for 2+1 dimensional gauge theories reads
@@
where $\omega=\frac{(2k_{0}+1)\pi}{N_{t}}$ with $k_{0}=0,\ldots,N_{t}-1$ are the Matsubara frequencies of the system, and $\vec{p}=(p_{1},p_{2})=\frac{2\pi\vec{k}}{N_{s}}$ with $k_{1,2}=0,\ldots,N_{s}-1$ are the allowed spatial momenta (in lattice units). The positive spectrum is then entirely contained in the finite interval $[\lambda_{(0)},\lambda_{(1)}]$ with $\lambda_{(0)}=\sin\omega_{0}$ and $\lambda_{(1)}=\sqrt{\max_{k}(\sin\omega_{k})^{2}+2}$. We will refer to this region as the bulk of the spectrum, and to modes in the regions $0\leq\lambda<\lambda_{(0)}$ and $\lambda_{(1)}<\lambda\leq 3$ as low and high modes, respectively.
@@
where $O_{l}$ denotes any eigenmode observable evaluated for mode $l$, and we have made explicit the dependence of the average on the spatial size $N_{s}$ of the system. The numerator in Eq. (9) is the spectral density, i.e., the local density of modes, that grows proportionally to the lattice volume. The suitably normalized spectral density,
@@ II.3 Center vortices
One can then ask if in typical gauge configurations one finds an infinite center vortex, i.e., a cluster whose size grows linearly with the lattice volume. This is the percolation problem we are interested in. As argued in Ref. [39], the presence of such a cluster would lead to a finite string tension and so to a confining theory. The cluster size can be measured by the number of links, or alternatively by the number of sites it contains. We choose the latter, and denote with $\mathcal{S}$ the fraction of dual sites occupied by the largest cluster in a given gauge configuration, i.e.,
@@
This quantity sets the concentration of active bonds in the corresponding percolation problem. Notice that the use of periodic boundary conditions in space and the topological constraints on the bond configurations make it a different problem from usual percolation in two dimensions.
@@ II.4 Eigenmode localization
The localization properties of modes in a given spectral region can then be determined by comparing the spectral statistics of the unfolded spectrum of the model, computed locally in the spectrum, with the theoretical predictions: As the system size increases, these observables will tend to the Poisson or RMT theoretical expectation depending on whether the modes are localized or delocalized in the spectral region under consideration. A point in the spectrum separating localized and delocalized modes is known as mobility edge and usually denoted $\lambda_{c}$. At a mobility edge the localization length $\xi$ diverges, typically as a power law, $\xi\sim|\lambda-\lambda_{c}|^{-\nu}$, and the system displays a second-order phase transition along the spectrum known as Anderson transition [25, 26, 27, 28, 29]. This implies that at the mobility edge the spectral statistics are volume-independent. Moreover, one can precisely determine the mobility edge, the critical statistics, and the critical exponent $\nu$ by means of a finite-size-scaling analysis of spectral statistics [44].
In the case at hand, the random matrix of interest is $H=-iD^{\rm stag}$, which is invariant under the anti-unitary transformation $T=\varepsilon K$, where $K$ denotes complex conjugation and $\varepsilon$ is defined in Eq. (7). Since $T^{2}=1$, $H$ belongs to the orthogonal class. [^2] The random matrix of interest satisfies also the chiral property $\{\varepsilon,H\}=0$ with $[\varepsilon,T]=0$, which puts it in the chiral orthogonal class for what concerns the statistical behavior of eigenvalues near zero. This distinction does not, however, affect the statistical properties in the bulk of the spectrum. The corresponding unfolded level spacing distribution can be approximated quite accurately with the so-called orthogonal Wigner surmise, $p_{\rm WS}$ [42]. An even better approximation is obtained using the family of distributions $p_{\rm B}(s,\omega)$, $\omega\in[0,1]$, known as Brody distributions [45],
@@
$\overline{I}_{s_{0}}(\lambda,N_{s})=\frac{\langle\sum_{l}\delta(\lambda-\lambda_{l})\theta(s_{0}-s_{l})\rangle}{\langle\sum_{l}\delta(\lambda-\lambda_{l})\rangle}\,.$ (25)
@@
Moreover, all points in the spectrum beyond the mobility edge are critical, with volume-independent spectral statistical properties. Indications of this behavior for the $\mathbb{Z}_{2}$ model were found in Ref. [19]. Similar indications were found also in 2+1 dimensional pure $\mathbb{Z}_{3}$ gauge theory in Ref. [20]. The divergence of the localization length shown in Eq. (26) leads one to expect the following scaling behavior for $I_{s_{0}}$ near the mobility edge,

### v2

# Center vortices and localized Dirac modes in the deconfined phase of (2+1)-dimensional lattice $\mathbb{Z}_{2}$ gauge theory

@@ Abstract
We study the deconfinement transition in (2+1)-dimensional lattice $\mathbb{Z}_{2}$ gauge theory both as a percolation transition of center vortices and as a localization transition for the low-lying Dirac modes. We study in detail the critical properties of the Anderson transition in the Dirac spectrum in the deconfined phase, showing that it is of BKT type; and the critical properties of the center-vortex percolation transition, showing that they differ from those of ordinary two-dimensional percolation. We then study the relation between localized modes and center vortices in the deconfined phase, identifying the simple center-vortex structures that mainly support the localized Dirac modes. As the system transitions to the confined phase, center vortices merge together into an infinite cluster, causing the low Dirac modes to delocalize.
@@ I Introduction
A growing amount of evidence shows that the localization properties of low-lying Dirac modes in a gauge theory are strongly connected with its confining properties: Low Dirac modes, while extended in the confined phase of a gauge theory, become localized in the deconfined phase, doing so exactly at the critical temperature if the phase transition is a genuine thermodynamic transition, and within the crossover region otherwise [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] (see Ref. [24] for a recent review). More precisely, in the deconfined phase the localized low Dirac modes are separated from the delocalized bulk modes by a critical point in the spectrum known as “mobility edge”. At the mobility edge the localization length diverges and a phase transition along the spectrum, known as Anderson transition [25, 26, 27, 28, 29], takes place. In the cases investigated so far, as the temperature decreases the mobility edge is pushed towards the origin of the spectrum, and localized modes disappear from the spectrum (possibly abruptly [17, 16]) as one crosses over to the confined phase. While the existence of a relation between localization and deconfinement is by now well established, the nature of this relation has not been fully elucidated yet.
@@
The connection between deconfinement and localization of low Dirac modes in gauge theories is qualitatively understood in terms of the “sea/islands” picture proposed in Ref. [5], further developed in Refs. [30, 31, 32, 24], and refined in Ref. [20]. In the deconfined phase, in the “physical sector” where Polyakov loops order along 1, selected by dynamical fermions or, in the case of pure gauge theories, by (infinitely) heavy external fermionic probes, a pseudogap opens in the Dirac spectrum due to the ordering of the Polyakov loop. In the resulting “sea” of ordered local Polyakov loops, gauge-field fluctuations that reduce the correlation across time slices, such as Polyakov-loop fluctuations away from 1, are localized on “islands”, typically well-separated in space. These fluctuations reduce the ability of Dirac modes to diffuse in space, and can “trap” the Dirac modes if these develop an eigenvalue below the gap, localizing the corresponding eigenvectors. This mechanism applies in the deconfined phase of a generic gauge theory, explaining the observed universal nature of the connection between localization and deconfinement; and, more generally, in phases where the Polyakov loop gets ordered, e.g., in the Higgs phase of the $\mathrm{SU}(2)$ Higgs model [22].
@@
The connection between localization and deconfinement shows up even in the simplest gauge theory with a deconfining transition, namely pure $\mathbb{Z}_{2}$ gauge theory in (2+1) dimensions [19]. This model is dual to an anisotropic three-dimensional Ising model [33] (summed over boundary conditions: see Refs. [34, 35] and references therein, and Ref. [20] for a detailed derivation), which allows for an accurate determination of the critical value of the gauge coupling [34] through the use of a cluster algorithm [36, 37]. While localization of low Dirac modes and its appearance at deconfinement were clearly demonstrated in Ref. [19], a precise determination of the mobility edge and of its temperature dependence, as well as a detailed study of the corresponding Anderson transition, were not undertaken.
@@
Gauge configurations in an Abelian model like the $\mathbb{Z}_{2}$ gauge theory are entirely determined by the simplest plaquettes and by the Polyakov loops in the various directions (with spatial Polyakov loops playing no important dynamical role in the large-volume limit). In the $\mathbb{Z}_{2}$ gauge theory, plaquette configurations on the direct lattice are in one-to-one correspondence with configurations of closed loops on the dual lattice, constructed by associating an “active” dual link to each nontrivial plaquette. Dual loops can further intersect each other, and are conveniently grouped into disconnected clusters of various sizes. Since the gauge group is Abelian and therefore equal to its own center, these clusters are trivially the same as the center vortices [38] of this model, and so we will refer to them also as center vortices in the following. Center vortices can be constructed for more general gauge groups as well, and have been proposed as the origin of confinement in gauge theories in Ref. [39]. The center-vortex picture of confinement has been the subject of intense research activity [40, 41, 42, 43, 44, 45, 46, 47, 48] that is still ongoing (see, e.g., Refs. [49, 50, 51, 52, 53]). We could do no justice to the vast related literature: we refer the interested reader to Ref. [38] for an extensive discussion of this picture, and to Refs. [49, 50, 51, 52, 53] for recent developments and updated lists of references.
@@
The center-vortex picture of confinement is particularly simple in the case of $\mathbb{Z}_{2}$ gauge theory, where the sum of the sizes of the center-vortex clusters equals the number of negative plaquettes of the gauge configuration on the direct lattice, and one can exactly recast the partition function in terms of dual clusters. The dynamics of $\mathbb{Z}_{2}$ gauge fields is then the dynamics of center vortices. In this language, confinement can be understood as center-vortex percolation, i.e., as the formation of a percolating cluster of dual active links, that leads to a nonvanishing string tension. Conversely, deconfinement corresponds to the disappearance of the percolating cluster, with only clusters of finite size surviving the large-volume limit. [^1] Notice that since a plaquette configuration is left unchanged by a center transformation, the center sector selected by the Polyakov loop, as it gets ordered and breaks center symmetry, plays no role in this dynamics. The idea of looking at the confinement-deconfinement transition as a center-vortex percolation transition was proposed in Refs. [42, 43], where finite-temperature pure $\mathrm{SU}(2)$ gauge theory in (3+1) dimensions was investigated. Reference [54] studied instead zero-temperature $\mathbb{Z}_{2}$ gauge theory in (2+1) dimensions, showing the presence of an infinite center-cluster in typical gauge configurations in the confined phase, and demonstrating its crucial role in establishing confinement. On the other hand, the behavior of clusters across the (bulk) deconfinement transition [55, 56] was not investigated. This was done more recently and in a different context in Ref. [57], that reports on a detailed study of percolation and its relation with the thermodynamic transition.
@@
Pure $\mathbb{Z}_{2}$ gauge theory in (2+1) dimensions is clearly the easiest setup in which one can study the nature of the gauge-field fluctuations relevant to localization, and how these relate to the confining properties of the theory. The connection between deconfinement and localization on the one hand, and that between confinement and percolation on the other, suggest that the nonpercolating center clusters expected in the deconfined phase should play an important role for localization. In this paper we show that this is indeed the case. Of course, establishing a connection between localized Dirac modes and nonpercolating center clusters in the present model has no direct bearing on the role played by vortices in the confinement mechanism in the physically relevant case of QCD. Nonetheless, it shows that at least in this model the features of gauge-field configurations responsible for the confinement-deconfinement transition are exactly the same responsible for the corresponding change in the localization properties of Dirac modes. Identifying the gauge-field structures that drive this change across the deconfinement transition in QCD and in other gauge theories could then be of guidance in the search for the correct microscopic mechanism behind confinement in realistic theories.
The relation between center vortices and the Dirac spectrum has been studied in pure $\mathrm{SU}(2)$ gauge theory in (3+1) dimensions both at zero [45, 47] and nonzero temperature [46, 47, 48], making use of the overlap discretization of the Dirac operator [58] (as well as of asqtad fermions [59] in Ref. [47]). As the main focus of those studies is the spectral density of the modes (although localization properties were also investigated in Refs. [47, 48]) and its connection with chiral symmetry breaking, using overlap fermions makes a difference due to their superior chiral properties. Here we are interested instead in the localization properties of Dirac modes and their relation with confinement, and for this kind of study the use of a discretization with good chiral properties is not required. Indeed, localization of low Dirac modes in the deconfined phase has been observed with a variety of fermion discretizations, with the same picture resulting from the use of chiral and nonchiral discretizations (see Ref. [24] for a review). A particularly clear example is pure $\mathrm{SU}(3)$ gauge theory in (3+1) dimensions, where localized modes appear exactly at the first-order deconfinement transition both in the staggered [12] and in the overlap spectrum [15, 17].
@@
In this paper we investigate in detail the localization of Dirac modes, the percolation of center vortices, and how they affect each other in finite-temperature pure $\mathbb{Z}_{2}$ gauge theory in (2+1) dimensions on the lattice, probed with external staggered fermions. In Sec. II, after briefly introducing the model in Sec. II.1 and staggered fermions in Sec. II.2, we discuss a few generalities about center vortices in Sec. II.3, and about localization and the sea/islands picture in Secs. II.4 and II.5. In Sec. III we present our numerical results. We first determine the nature of the Anderson transition and the temperature dependence of the mobility edge in Sec. III.1. We then study the percolation of center vortices across the deconfinement transition in Sec. III.2 and the correlation between center clusters and localized modes in Sec. III.3, and test the refined sea/islands picture of Ref. [20] in Sec. III.4. We then identify the features of center vortices mostly relevant to localization in Sec. III.5. Finally, in Sec. IV we present our conclusions and discuss prospects for future studies. A brief discussion of some unexplained features of the staggered spectrum noted in Ref. [19] is reported in Appendix A.
@@ II Localization and center vortices in $\mathbb{Z}_{2}$ lattice gauge theory at finite temperature
In this section, after briefly describing (2+1)-dimensional pure $\mathbb{Z}_{2}$ gauge theory on the lattice and how to probe it with external staggered fermions, we summarize the relevant aspects of center vortices and their percolation, and of Dirac mode localization.
II.1 Finite-temperature $\mathbb{Z}_{2}$ lattice gauge theory in (2+1) dimensions
We consider pure $\mathbb{Z}_{2}$ gauge theory on a (2+1)-dimensional hypercubic $N_{t}\times N_{s}^{2}$ lattice. Lattice sites are denoted by $n=(n_{1},n_{2},n_{3})=(t,\vec{x})$, where $10\leq t=n_{1}\leq N_{t}-$ 1, $10\leq x_{1,2}=n_{2,3}\leq N_{s}-$ 1. The dynamical variables of the model are the link variables $U_{\mu}(n)=\pm 1$, taking values in $\mathbb{Z}_{2}$, and associated with the oriented lattice links connecting a site $n$ with its neighbors $n+\hat{\mu}$, where $\mu=1,2,3$ and $\hat{\mu}$ is the unit lattice vector in direction $\mu$. Periodic boundary conditions are imposed in all directions.
@@ II.1 Finite-temperature $\mathbb{Z}_{2}$ lattice gauge theory in (2+1) dimensions
Together with the analogs of $P(\vec{x})$ winding around the spatial directions, these quantities determine the link configuration up to a gauge transformation.
@@ II.2 Staggered Dirac spectrum
We probe the gauge-field configurations using external fermion fields by studying the spectrum of the staggered Dirac operator, which for (2+1)-dimensional gauge theories reads
@@
where $\omega=\frac{(2k_{0}+1)\pi}{N_{t}}$ with $k_{0}=0,\ldots,N_{t}-1$ are the Matsubara frequencies of the system, and $\vec{p}=(p_{1},p_{2})=\frac{2\pi\vec{k}}{N_{s}}$ with $k_{1,2}=0,\ldots,N_{s}-1$ are the allowed spatial momenta (in lattice units). The positive spectrum is then entirely contained in the finite interval $[\lambda_{(0)},\lambda_{(1)}]$ with $\lambda_{(0)}=\sin\frac{\pi}{N_{t}}$ the lowest lattice Matsubara frequency, and $\lambda_{(1)}=\sqrt{\max_{\omega}(\sin\omega)^{2}+2}$. We will refer to this region as the bulk of the spectrum, and to modes in the regions $0\leq\lambda<\lambda_{(0)}$ and $\lambda_{(1)}<\lambda\leq 3$ as low and high modes, respectively.
@@
where $O_{l}$ denotes any eigenmode observable evaluated for mode $l$, and we have made explicit the dependence of the average on the spatial size $N_{s}$ of the system. The denominator in Eq. (9) is the spectral density, i.e., the local density of modes, that grows proportionally to the lattice volume. The suitably normalized spectral density,
@@ II.3 Center vortices
One can then ask if in typical gauge configurations one finds an infinite center vortex, i.e., a cluster whose size grows linearly with the lattice volume. This is the percolation problem we are interested in. As argued in Ref. [43], the presence of such a cluster would lead to a finite string tension and so to a confining theory. That this happens in the $\mathbb{Z}_{2}$ gauge theory was shown in Ref. [54]. The cluster size can be measured by the number of links, or alternatively by the number of sites it contains. We choose the latter, and denote with $\mathcal{S}$ the fraction of dual sites occupied by the largest cluster in a given gauge configuration, i.e.,
@@
This quantity sets the concentration of active bonds in the corresponding percolation problem. Notice that the use of periodic boundary conditions in space and the topological constraints on the bond configurations, as well as the presence of an additional (but finite) dimension, make it a different problem from usual percolation in two dimensions (see Refs. [60, 61, 62, 63]).
@@ II.4 Eigenmode localization
The localization properties of modes in a given spectral region can then be determined by comparing the spectral statistics of the unfolded spectrum of the model, computed locally in the spectrum, with the theoretical predictions: As the system size increases, these observables will tend to the Poisson or RMT theoretical expectation depending on whether the modes are localized or delocalized in the spectral region under consideration. A point in the spectrum separating localized and delocalized modes is known as mobility edge and usually denoted $\lambda_{c}$. At a mobility edge the localization length $\xi$ diverges, typically as a power law, $\xi\sim|\lambda-\lambda_{c}|^{-\nu}$, and the system displays a second-order phase transition along the spectrum known as Anderson transition [25, 26, 27, 28, 29]. This implies that at the mobility edge the spectral statistics are volume independent. Moreover, one can precisely determine the mobility edge, the critical statistics, and the critical exponent $\nu$ by means of a finite-size-scaling analysis of spectral statistics [67].
In the case at hand, the random matrix of interest is $H=-iD^{\rm stag}$, which is invariant under the antiunitary transformation $T=\varepsilon K$, where $K$ denotes complex conjugation and $\varepsilon$ is defined in Eq. (7). Since $T^{2}=1$, $H$ belongs to the orthogonal class. [^2] The random matrix of interest satisfies also the chiral property $\{\varepsilon,H\}=0$ with $[\varepsilon,T]=0$, which puts it in the chiral orthogonal class for what concerns the statistical behavior of eigenvalues near zero. This distinction does not, however, affect the statistical properties in the bulk of the spectrum. The corresponding unfolded level spacing distribution can be approximated quite accurately with the so-called orthogonal Wigner surmise, $p_{\rm WS}$ [65]. An even better approximation is obtained using the family of distributions $p_{\rm B}(s,\omega)$, $\omega\in[0,1]$, known as Brody distributions [68],
@@
$\overline{I}_{s_{0}}(\lambda,N_{s})=\frac{\langle\sum_{l}\delta(\lambda-\lambda_{l})\theta(s_{0}-s_{l})\rangle}{\langle\sum_{l}\delta(\lambda-\lambda_{l})\rangle}\,,$ (25)
where $\theta(x)$ is the Heaviside step function.
@@
Moreover, all points in the spectrum beyond the mobility edge are critical, with volume-independent spectral statistical properties. Indications of this behavior for the $\mathbb{Z}_{2}$ model were found in Ref. [19]. Similar indications were found also in (2+1)-dimensional pure $\mathbb{Z}_{3}$ gauge theory in Ref. [20]. The divergence of the localization length shown in Eq. (26) leads one to expect the following scaling behavior for $I_{s_{0}}$ near the mobility edge,